OpenStudy (anonymous):

When a mole of lithium reacts chemically with hydrogen to form lithium hydride, the energy released is found to be roughly 90 joules. How much mass is lost? Answer 1.0 x 10-15 kg 2.7 x 1010 kg 1.0 x 1015 kg 3.0 x 10-7 kg

OpenStudy (aaronq):

they want you to use \(\sf E=mc^2\)

OpenStudy (anonymous):

E=mc^2 --> m= E/c^2 = 90 J / (3*10^8)^2 =1.0x10^-15 kg